# Quality-Driven Curation of Remote Sensing Vision-Language Data via Learned Scoring Models

### v1

@@ 5 Conclusion
Vision-language data serves as the fundamental building block for training VLMs. However, the quality control and data curation for RS-specific VLMs have not been fully addressed. In this study, we explore the development of parameter-wise scoring models for high-quality data selection. Through careful construction of RS preference datasets and the training of our ScoreRS model, we demonstrate that current vision-language datasets are far from optimal. Our findings show that using just 30% of quality-filtered data achieves superior performance compared to the complete dataset in both CLIP training and large VLM fintuning. Furthermore, we investigate the application of ScoreRS in RL training and BoN selections. Both applications demonstrate that ScoreRS can enhance VLM capabilities in solving complex and challenging tasks. We anticipate that this study will encourage the RS community to place more emphasis on vision-language data quality control and the strategic utilization of high-quality data.
Appendix
@@
AID METERML NWPU SIRI-WHU WHU-RS19 Avg. HR-Comp. HR-Pres. LR-Comp. LR-Pres. LR-R-U Avg. VG-DOIR LHRS-Bench
@@
6 Additional Result
6.1 CLIP Training
The results, presented in Tab. 7, further demonstrate that filtering with our ScoreRS yields superior performance compared to both using the complete dataset and applying CLIP-score filtering. We observe that directly fine-tuning CLIP with the entire Skyscript dataset actually resulted in performance inferior to the original CLIP model. We attribute this to the rule-based caption construction method used in Skyscript, which introduces significant redundancy and thus necessitates more careful data selection for effective fine-tuning.
6.2 PPO Training
@@
We explored using our ScoreRS directly as a reward model for RL through proximal policy optimization (PPO) [32]. We employ our finetuned Qwen2VL-7B-RS as the policy model, while the critic model is initialized with the same Qwen2VL-7B-RS architecture but with its language head replaced by a learnable linear layer. We utilize the same dataset described in Sec. 3.2 for this training process. LoRA is applied to all linear layers in the policy model with rank and $\alpha$ parameters both set to 64. For text generation, we configure a sampling temperature of 0.95 and top-p of 0.9, with maximum new token generation limited to 512 tokens. The implementation uses a rollout batch size of 16, with PPO buffer size of 8 and 4 PPO epochs. We initialize the KL penalty coefficient at 2 and gradually increase it to 6 throughout the training process. The $\lambda$ parameter for Generalized Advantage Estimation (GAE) [31] is set to 0.95. For optimization, we use a learning rate of $1\times 10^{-5}$ with a 0.1 warmup ratio and cosine learning rate decay schedule.
@@
The results presented in Tab. 8 reveal that the model after PPO training performs inferiorly to its base variant across almost all evaluation datasets. We suspect this underperformance stems from suboptimal hyperparameter selection and reward calculation. Since we directly use our ScoreRS as reward model without any normalization or reference-based approach, this may be caused by reward hacking [1]. Moreover, given the considerable number of parameters associated with the PPO algorithm, we expect that developing comprehensive reward calculation strategies and conducting thorough hyperparameter searches remain important directions for future work.
7 Preference Data Generation
7.1 Image-Caption Preference Dataset
Caption Generation Setting
Caption Judgment Setting
For the selection of captions in our preference pairwise data, we employ GPT-4o (GPT-4o-2024-05-13). The complete prompt used for this judgment is provided in Tab. 10.
Data Samples
We provide several representative samples from our image-caption preference dataset in Fig. 5 to illustrate the quality of the collected data.
7.2 Vision Instruction Preference Dataset
Judgment Setting
We provide the prompt in Tab. 11 for using Qwen2VL-72B as judger to select the better answer for the input question.
RS Specific Category and Question
We manually design RS specific questions across 12 categories: basic visual recognition, spatial analysis, environmental assessment, urban analysis, agricultural analysis, disaster assessment, geological features, infrastructure analysis, temporal understanding, advanced reasoning, quantitative assessment, and color spectral analysis. These domain-specific questions are presented in Tab. 12 and Tab. 13.
Generation Prompt
We utilized LHRS-Bot-Nova and Qwen2VL-7B to answer the RS-specific questions. The prompt template employed for this question-answering task is presented in Tab. 14.
Rephrase Prompt
We employ Qwen2.5-13B to rephrase the manually designed questions to enhance question diversity. The prompt used for this rephrasing process is presented in Tab. 15.
Data Samples
We provide serveal samples from our vision instruction preference dataset in Fig. 6.
8 Experimental Setting
8.1 ScoreRS Training

### v2

@@ 5 Conclusion
Vision-language data serves as the fundamental building block for training VLMs. However, the quality control and data curation for RS-specific VLMs have not been fully addressed. In this study, we explore the development of parameter-wise scoring models for high-quality data selection. Through careful construction of RS preference datasets and the training of our ScoreRS model, we demonstrate that current vision-language datasets are far from optimal. Our findings show that using just 30% of quality-filtered data achieves superior performance compared to the complete dataset in both CLIP training and large VLM finetuning. Furthermore, we investigate the application of ScoreRS in RL training and BoN selections. Both applications demonstrate that ScoreRS can enhance VLM capabilities in solving complex and challenging tasks. We anticipate that this study will encourage the RS community to place more emphasis on vision-language data quality control and the strategic utilization of high-quality data.
6 Acknowledgement
We would like to thank Haoqin Tu, Zhuo Zheng, and Mengke Zhu for their valuable discussions and feedback.
Appendices
Appendix A Quality Dimension and Scoring System
A.1 Quality Dimension
The detailed descriptions and clarifications for each dimension are provided below:
Relevance: The text accurately describes the content visible in the remote sensing image. It does not introduce objects, features, or relationships that are not present (i.e., avoids hallucination). All key elements mentioned in the text are verifiable in the image.
Specificity & Detail: The text provides specific and detailed information rather than being vague or overly general. It names identifiable objects, describes their characteristics (e.g., size, shape, color if discernible and relevant), and their spatial relationships.
Completeness: The text covers the most salient and important features of the image relevant to common RS tasks (e.g., land use classification, object detection, change detection). It doesn’t omit obvious, critical elements. For question-answering, the answer should address the question if the information is present.
Clarify & Fluency: The text is grammatically correct, well-structured, and easily understandable. It uses clear, complete, and unambiguous language.
Semantic Richness: The text captures semantic information that is particularly useful for downstream RS applications. This could include types of land cover (e.g., "deciduous forest," "industrial zone," "low-density residential"), specific object classes ("solar farm," "roundabout," "cooling towers"), or activities ("active construction site").
Distinction Between Specificity & Detail, Completeness, and Semantic Richness:
We delineated three potentially overlapping quality dimensions to ensure non-redundant evaluation:
(1) Specificity & detail focuses on the level of precise information provided (e.g., "high-rise commercial buildings" vs. just "buildings"). It’s about granularity and precision in what is described.
(2) Completeness focuses on coverage of all important elements in the image. It’s about breadth - whether all significant features are mentioned, regardless of how detailed each mention is.
(3) Semantic Richness focuses on the use of domain-specific terminology and concepts relevant to remote sensing (e.g., "center-pivot irrigation" vs. "circular fields").
These dimensions can vary independently—a description might be highly specific yet incomplete, complete yet lacking domain terminology, or semantically rich yet insufficiently detailed—justifying their separate evaluation.
A.2 Scoring System
We provide our detailed scoring criteria for image-caption pairs in Table 16 and for instruction samples in Tables 17 and 18.
Appendix B Additional Detail for Preference Data Generation
B.1 Deduplication
Image-Caption:
We implement a rigorous image deduplication process for the LHRS-Align [39] to ensure image representativeness for building our image-caption preference dataset: (1) Feature extraction using the SSCD copy detection model [45, 18] to compute image embeddings. (2) Similarity computation through cosine distance computation between image embeddings, with a carefully tuned threshold of 0.65 (empirically determined through experiments across the range 0.6-0.9). (3) Duplicate grouping via connected-components algorithm, preserving one image-text pair per component. Finally, this deduplication pipeline yields 76K distinctive representative RS images.
Vision Instruction:
We employ a two-stage similarity-based filtering process to ensure diversity of instruction sample: (1) Text similarity: let $Q=\{T_{1},T_{2},\ldots,T_{n}\}$ be the set of all questions extracted from the conversations. We compute their semantic embeddings using the BGE [6] model, where for any two queries $T_{i},T_{j}\in Q:\text{sim}_{\text{text}}(T_{i},T_{j})=\cos(\text{BGE}(T_{i}),\text{BGE}(T_{j}))$; (2) Image similarity: For query pairs where $\text{sim}_{\text{text}}(T_{i},T_{j})>0.65$, we further examine their corresponding images $I_{i},I_{j}$ using SSCD [45] embeddings: $\text{sim}_{\text{image}}(I_{i},I_{j})=\cos(\text{SSCD}(I_{i}),\text{SSCD}(I_{j}))$. When both $\text{sim}_{\text{text}}(T_{i},T_{j})$ and $\text{sim}_{\text{image}}(I_{i},I_{j})$ exceed 0.65, we retain only one question-image sample from the pair to minimize redundancy.
B.2 Model Selection
For the preference generation policy models, we aimed to maximize diversity. However, considering the high costs associated with adding multiple models (including generation and evaluation costs), we selected 3 policy models for image-caption preference dataset generation and 2 for vision instruction preference dataset creation.
Regarding model families, we ensured inclusion of two model types: RS-specific VLMs and general VLMs, to increase answer diversity. For example, RS-specific VLMs typically include domain-specific terminology such as "high resolution, true color", while general VLMs often add contextual descriptors like "aerial, satellite".
For RS-specific VLMs, we selected LHRS-Bot-Nova [30] based on its superior performance in our preliminary evaluations. For general VLMs, we aim to select the highest-performing available models. At the time of dataset preparation, the best open-source VLMs are Qwen2VL and InternVL2.5, so we selected these model families using the same parameter scale (7B) as LHRS-Bot-Nova for policy generation.
We acknowledge that model size, model family, prompting strategies, and generation hyperparameters all influence the final preference dataset [29, 23], and synthetic data generation itself represents a critical research topic [41, 2]. In this work, we focused primarily on a reasonable model selection strategy while emphasizing diversity and domain specificity of the RS images and instructions used to generate preference data. We leave for future work the exploration of how different policy models affect preference data quality and scoring model performance.
B.3 Generation Parameters and Prompts
B.3.1 Image-Caption Preference Dataset
Caption Generation Setting
Caption Judgment Setting
For the selection of captions in our preference pairwise data, we employ GPT-4o (GPT-4o-2024-05-13). The complete prompt used for this judgment is provided in Table 19.
Data Samples
We provide several representative samples from our image-caption preference dataset in Figure 10 to illustrate the quality of the collected data.
B.3.2 Vision Instruction Preference Dataset
Judgment Setting
We provide the prompt in Table 20 for using Qwen2VL-72B as judger to select the better answer for the input question.
RS Specific Category and Question
We manually design RS specific questions across 12 categories: basic visual recognition, spatial analysis, environmental assessment, urban analysis, agricultural analysis, disaster assessment, geological features, infrastructure analysis, temporal understanding, advanced reasoning, quantitative assessment, and color spectral analysis. These domain-specific questions are presented in Table 21 and Table 22.
Generation Prompt
We utilized LHRS-Bot-Nova and Qwen2VL-7B to answer the RS-specific questions. The prompt template employed for this question-answering task is presented in Table 23.
Rephrase Prompt
We employ Qwen2.5-13B to rephrase the manually designed questions to enhance question diversity. The prompt used for this rephrasing process is presented in Table 24.
Data Samples
We provide serveal samples from our vision instruction preference dataset in Figure 11.
Table 7: Human validation of preference judgments. ’Model-Human Agreement’ measures the model’s accuracy against human judges. ’Inter-Annotator Agreement’ measures the consistency among human judges on a shared subset of samples.
Model-Human Agreement Inter-Annotator Agreement
Accuracy @ ICPD Accuracy @ VIPD Human Consistency @ ICPD Human Consistency @ VIPD
90.06% (1351/1500) 96.80% (1452/1500) 88% (44/50) 96% (48/50)
B.4 Further Human Validation of Preference Data
While powerful, using general-purpose models like GPT-4o and Qwen2VL-72B for preference judgments warrants a careful validation of their accuracy. We chose these models over existing domain-specific VLMs (which are typically smaller,  13B parameters) because they demonstrate superior instruction-following capabilities for the nuanced task of judging responses. For instance, our preliminary human validation showed that GPT-4o achieved 92.6% agreement, significantly outperforming a domain-specific model like LHRS-Bot-Nova (82.8%). Despite limited direct exposure to remote sensing data, the diverse training of GPT-4o and Qwen2VL-72B provides the robust, context-dependent reasoning required for this task.
To further validate the accuracy of our preference dataset, we conducted a dedicated human annotation experiment. We selected 3,000 samples of preference data, distinct from those in our main manuscript validation, comprising 1,500 from the image-caption preference dataset (ICPD) and 1,500 from the vision-instruction preference dataset (VIPD). Three human annotators were tasked with a binary judgment (Yes/No) on whether the chosen positive sample was qualitatively better than the negative sample. Additionally, to assess the reliability of our human evaluation, all three annotators evaluated a shared set of 100 common samples (50 from ICPD, 50 from VIPD) to establish inter-annotator agreement.
The results, summarized in Table 7, show that the judgments from our models align well with human preferences. The higher consistency and model alignment on VIPD tasks suggest that instruction-based responses are easier to evaluate than the more subjective, lengthy image captions, where human preferences can naturally vary. While this curation strategy proves effective, we acknowledge it is not perfect. Future work could incorporate more complex methods, such as Retrieval-Augmented Generation (RAG) with human-annotated data or self-reflection mechanisms, to further refine judgment quality.
Appendix C Additional Result
C.1 Score Model Training
Directly using the output from the final value head as a score creates challenges in defining the range, which affects not only the relative importance of data but also reward calculation in RL training (notably, algorithms like GRPO typically require a reward range of 0-1). To address this issue, we attempt to resolve it at the source by adding normalizing functions after the value head output.
We experiment with sigmoid, tanh, and softplus functions (though softplus only ensures positive scores). The training loss is presented in Figure 5. Despite trying different initializations for the final value head, we observe that using tanh and sigmoid as normalization functions led to easy convergence during training (which we interpret as the scores lying in the saturated region of these functions). Furthermore, since we observe that softplus performed worse than using no normalization at all (labeled as "raw" in Figure 5), we opt to use the latter setting during the score model training.
Figure 5: Training loss of score model with different normalization functions
C.2 Ablation Study
C.2.1 Influence of Model Size
We analyze the impact of initialized model size on the effectiveness of our scoring model. Considering the computational cost for full-parameter tuning of models larger than 7B, we evaluate the smaller Qwen2VL-2B model as a comparison to Qwen2VL-7B. For simplicity, we compare the results using the joint training strategy introduced in Section 4.4. To ensure a fair comparison, we use the same model family (Qwen2VL) and identical optimization steps (i.e., same batch size). We sweep through a wide range of learning rates from $1\times 10^{-7}$ to $1\times 10^{-4}$ for Qwen2VL-2B, and find that $8\times 10^{-5}$ yields the best results (according to the training loss), compared to $1\times 10^{-6}$ for Qwen2VL-7B.
Figure 6: Comparison of scoring models with different sizes on training loss and performance as BoN selector for LHRS-Bench
Table 8: Ablation analysis on the effectiveness of scoring model size for distinguishing between good and bad text sample on the hold-out preference set. ICPD = Image-Caption Preference Dataset, VIPD = Vision Instruction Preference Dataset
Accuracy @ ICPD Accuracy @ VIPD Accuracy
Jointly Training w. 7B 83.13 81.07 82.09
Jointly Training w. 2B 71.53 79.63 75.58
We compare the two variants’ training loss, accuracy on the held-out preference set as described in Section 4.4, and their performance as BoN selectors. From the training loss and BoN results shown in Figure 6, we observe that the smaller model converges at a higher loss compared to the larger model. On LHRS-Bench as a selector, the smaller selector performs inferior to the larger selector and tends to saturate when given more candidate samples, indicating that smaller models struggle to identify the best candidate among different samples. On the hold-out preference evaluation set presented in Table 8, we see that the 2B model performs approximately 7% worse compared to the 7B model, further demonstrating its reduced effectiveness in distinguishing good responses from bad ones compared to the larger model.
These results confirm that larger scoring models indeed provide better performance, which aligns with related work [17]. We anticipate that future work could develop even larger, more advanced scoring models to bring additional benefits to the RS community.
C.2.2 Influence of Model Initialization
Next, we tackle a critical question: whether initializing the scoring model with an RS-specific VLM instead of a general VLM (QwenVL2-7B in our case) provides benefits. We use our fine-tuned Qwen2VL-FT from Table 3 (i.e., the model fine-tuned with all VHM datasets) to initialize the scoring model and conduct the three-stage training with the same preference data. In this setting, we maintain the original base architecture (Qwen2VL) and training recipe to ensure a fair comparison. For evaluation, we use the different scoring models to filter the VHM vision instruction dataset with a fixed selection ratio of 60%, and use the filtered instruction dataset to fine-tune Qwen2VL-FT (ScoreRS-P30) from Table 3. The fine-tune setting is the same as Section D.4. We also evaluate them as BoN selectors on VG-DIOR and LHRS-Bench with our Qwen2VL-7B-RS-R1. The BoN evalution setting is the same as Section 4.4.
Figure 7: BoN comparison on different score model with different initialzietion base model. ScoreRS is initialized with Qwen2VL-7B, RS init. is initialized with Qwen2VL-FT from Table 3, and RS init. is initialized with Qwen2VL-FT (ScoreRS-P30S60) from Table 3
Table 9: Comparison of different initialization strategies for the scoring model. We initialize scoring models with different base models, use them to filter the VHM instruction dataset with a 60% selection ratio, and fine-tune Qwen2VL-FT (ScoreRS-P30) from Table 3 with the filtered data. ScoreRS is initialized with Qwen2VL-7B, RS init. is initialized with Qwen2VL-FT from Table 3, and RS init. is initialized with Qwen2VL-FT (ScoreRS-P30S60) from Table 3
RS Classification RSVQA Grounding General Knowledge
AID METERML NWPU SIRI-WHU WHU-RS19 Avg. HR-Comp. HR-Pres. LR-Comp. LR-Pres. LR-R-U Avg. VG-DIOR LHRS-Bench
Qwen2-VL-FT (ScoreRS) 85.66 74.86 89.49 73.33 92.80 83.23 82.00 68.90 89.68 86.63 87.00 82.84 55.58 66.58
Qwen2-VL-FT (RS init.) 82.43 73.52 88.60 74.56 92.40 82.30 83.20 67.20 86.33 91.40 82.00 82.03 53.47 66.29
Qwen2-VL-FT (RS init.) 84.90 75.38 89.20 74.55 94.60 83.73 83.40 68.90 88.87 90.96 86.00 83.63 56.21 66.90
The results are presented in Table 9 and Figure 7. From these results, we observe that initializing the scoring model with an RS-specific VLM yields inferior performance in both data filtering quality and BoN selection compared to using the general Qwen2VL as the base model. We hypothesize that this occurs because Qwen2VL-FT is fine-tuned with the complete VHM dataset, which likely contains low-quality vision-language data that biases the base model toward a domain that is difficult to optimize for high-quality sample selection.
To test this hypothesis, we initialize the scoring model with an RS-specific VLM that is fine-tuned with ScoreRS-filtered data—specifically, Qwen2VL-FT (ScoreRS-P30S60) from Table 3. With this modification, the new scoring model (RS init.) performs better on both data filtering and as a BoN selector, which supports our hypothesis.
Our analysis indicates that if we can ensure the RS-specific VLM is of high quality (though "high quality" may be difficult to precisely define, it can be approximated through data that has undergone cleaning through a qualified pipeline), specialized initialization of the scoring model will yield better performance.
C.3 CLIP Training
C.3.1 Evaluation on Skyscript Dataset
@@
The results, presented in Table 10, further demonstrate that filtering with our ScoreRS yields superior performance compared to both using the complete dataset and applying CLIP-score filtering. We observe that directly fine-tuning CLIP with the entire Skyscript dataset actually resulted in performance inferior to the original CLIP model. We attribute this to the rule-based caption construction method used in Skyscript, which introduces significant redundancy and thus necessitates more careful data selection for effective fine-tuning.
C.3.2 Evaluation on Extreme Data Filtering Scenarios
We investigate the impact of extreme data filtering by evaluating performance at various data saving ratios, as shown in Table 11. The results demonstrate that data quantity remains crucial for maintaining high performance, particularly for classification tasks. Extreme filtering ratios significantly degrade classification accuracy, likely because datasets with numerous classes require a sufficient number of samples per class to learn effectively. In contrast, retrieval tasks appear more resilient to data reduction. This suggests that even a small, high-quality subset of data can effectively capture the essential text-to-image alignment concepts required for retrieval, thus maintaining comparable performance even under extreme pruning.
Table 11: Performance of different data filtering methods at various saving ratios. Lower saving ratios indicate more extreme data filtering.
Saving Ratio Method Classification Retrieval
30% ScoreRS 70.97 27.11
CLIP-Score 69.98 26.36
20% ScoreRS 68.44 27.01
CLIP-Score 67.70 26.69
10% ScoreRS 63.19 26.16
CLIP-Score 62.89 26.05
Table 12: Performance comparison of different filtering methods on the curated RemoteCLIP dataset (30% saving ratio). The classification benchmark includes NWPU, AID, METERML, WHU-RS19, SIRI-WHU, fMoW, and EuroSAT. The retrieval benchmark uses UCM and RSICD.
Method Classification Avg @ 1 Retrieval Avg @ 1
ScoreRS 70.97 27.11
CLIP-Score 69.98 26.36
SigCLIP-2 68.24 26.47
CLIP-LAION-RS 70.59 26.41
C.3.3 Comparison with State-of-the-Art Filtering Methods
To further evaluate the effectivness of our ScoreRS, we conduct an analysis using robust, publicly available models as baselines. We noted that CLIP-LAION-RS [62], which was fine-tuned on a remote sensing (RS) subset of the well-curated LAION-2B dataset, outperforms standard CLIP-filtered versions. We, therefore, curated the RemoteCLIP dataset using scores from both SigCLIP-2 (L/16-256) and CLIP-LAION-RS (L/14), comparing their filtering performance against our ScoreRS method. For this experiment, we used a fixed data saving ratio of 30%.
The results are presented in Table 12. Although SigCLIP-2 generally exhibits strong performance on various tasks compared to the original CLIP models, factors such as data distribution and training recipes may lead to different outcomes in the specialized remote sensing domain. The SigCLIP-2 filtered version shows inferior performance on classification tasks, while the CLIP-LAION-RS filtered version shows a marked improvement over the standard CLIP-Score. Notably, our ScoreRS filter demonstrates the best performance among all baselines, achieving the highest scores in both classification and retrieval.
From these experiments, it is clear that ScoreRS is the first scoring model in the RS community capable of effectively evaluating diverse types of vision-language data. Additionally, thanks to the strong capabilities of its Qwen2VL backbone, it can be extended to evaluate complex multi-image vision-language data, such as change captioning tasks.
C.4 RL Comparison
C.4.1 Compare with DPO
We compare the effectiveness of using ScoreRS as a reward function for open-ended questions with GPRO against Direct Preference Optimization (DPO)[47]. Specifically, we use the 26K vision instruction preference pairs (excluding the RS-specific preference data) constructed in Section 3.2.2 for DPO training. Starting with our fine-tuned Qwen2VL-7B-RS model (Table 5), we conduct full-parameter DPO training using LLaMA-Factory and compare the results with Qwen2VL-7B-Zero (Table 5). For a fair comparison, we train two DPO variants: one using 8K preference pairs (comprising 4K open-ended questions and 4K closed-ended questions), and another using the full 26K preference dataset. We evaluate the resulting models on VG-DIOR and LHRS-Bench evaluation datasets.
The results are presented in Table 13. We observe that with equivalent data scale, DPO performs inferior to GPRO training, while increasing the data scale for DPO yielded performance on par with or slightly better than GPRO training. This observation aligns with the established conclusion that RL training is typically more data-efficient than SFT [5, 11] (here, we consider DPO as a generalized form of SFT). In our case, the data efficiency advantage is approximately 3× (8K vs. 26K).
Figure 8: Comparison with GPT-4o as reward function for open-ended questions
Table 13: Comparison between ScoreRS-based GPRO and DPO RL training. DPO/8K represents training with the same data scale as GPRO training, while DPO/26K represents training with the complete preference dataset for DPO training
VG-DIOR LHRS-Bench
Qwen2VL-7B-RS-DPO/8K 58.97 66.45
Qwen2VL-7B-RS-DPO/26K 59.71 67.19
Qwen2VL-7B-RS-Zero 59.64 67.21
C.4.2 Compare with GPT-4o as Reward Model
Since our preference dataset is largely built using GPT-4o as a judge (excluding closed-ended questions) based on our scoring system, a natural question arises: would using the same scoring system with GPT-4o as the reward model for calculating rewards for open-ended questions be more effective? To investigate this, we utilize GPT-4o for reward calculation based on our scoring system and normalized the total reward (e.g., average reward / 5) as the final reward for open-ended questions. We start with the Qwen2VL-7B-RS model from Table 5 and use the same training dataset and experimental settings as described in Section 4.2. We compare these results with those obtained using our ScoreRS(i.e., compared with Qwen2VL-7B-RS-Zero).
We plot the total reward, response length, and evaluation results on LHRS-Bench during the training process in Figure 8. The results show that despite our dataset being largely built with GPT-4o, using GPT-4o as the open-ended reward model performed inferior to using our ScoreRS. We suspect two main reasons for this outcome. First, since our ScoreRS is fine-tuned with large-scale RS-specific preference data, although the dataset is built using GPT-4o judgments, it acquires substantial domain-specific knowledge (or simply put, developed a bias toward this specific domain) that helps it better distinguish between good and better responses. Second, the closed-ended preference dataset enables specific capabilities in ScoreRS. Moreover, we should note that using GPT-4o as a reward model during the training process is significantly more costly than using ScoreRS  (here, cost refers to money, with the single training run costing approximately \$360).
C.5 BoN Selection
We further evaluate our ScoreRS on more advanced VLMs, including InternVL3-8B [55] and Qwen2.5VL-7B [56]. To validate the effectiveness of our approach, we compare it with majority voting (i.e., selecting the most common candidate as the final answer). The generation sampling parameters are the same as those described in Section 4.3. The results are presented in Figure 9.
Our findings clearly demonstrate that even with more advanced (i.e., more optimized) models, our ScoreRS can boost performance as a BoN selector and outperform the majority voting selection mechanism. Notably, for the RS vision grounding task, since the bounding boxes generated vary across different candidate samples, majority voting often fails to improve model performance, while our ScoreRS consistently delivers improvements on this task.
Figure 9: BoN comparison on more advanced models with ScoreRS as selector compared to majority voting
@@ C.5 BoN Selection
AID METERML NWPU SIRI-WHU WHU-RS19 Avg. HR-Comp. HR-Pres. LR-Comp. LR-Pres. LR-R-U Avg. VG-DIOR LHRS-Bench
@@
C.6 PPO Training
@@
We explored using our ScoreRS directly as a reward model for RL through proximal policy optimization (PPO) [50]. We employ our finetuned Qwen2VL-7B-RS as the policy model, while the critic model is initialized with the same Qwen2VL-7B-RS architecture but with its language head replaced by a learnable linear layer. We utilize the same dataset described in Section 4.2 for this training process. LoRA is applied to all linear layers in the policy model with rank and $\alpha$ parameters both set to 64. For text generation, we configure a sampling temperature of 0.95 and top-p of 0.9, with maximum new token generation limited to 512 tokens. The implementation uses a rollout batch size of 16, with PPO buffer size of 8 and 4 PPO epochs. We initialize the KL penalty coefficient at 2 and gradually increase it to 6 throughout the training process. The $\lambda$ parameter for Generalized Advantage Estimation (GAE) [49] is set to 0.95. For optimization, we use a learning rate of $1\times 10^{-5}$ with a 0.1 warmup ratio and cosine learning rate decay schedule.
@@
The results presented in Table 14 reveal that the model after PPO training performs inferiorly to its base variant across almost all evaluation datasets. We suspect this underperformance stems from suboptimal hyperparameter selection and reward calculation. Since we directly use our ScoreRS as reward model without any normalization or reference-based approach, this may be caused by reward hacking [1]. Moreover, given the considerable number of parameters associated with the PPO algorithm, we expect that developing comprehensive reward calculation strategies and conducting thorough hyperparameter searches remain important directions for future work.
Appendix D Experimental Setting
D.1 Hardware and Framework
All our experiments are conducted on 2 nodes, each equipped with 8 $\times$ A100-80G GPUs. For training the scoring model, we develope a customized training framework based on the OLMo training framework [^2] https://github.com/allenai/OLMo. For fine-tuning the large VLMs, we utilized the LLaMA-Factory training framework [^3] https://github.com/hiyouga/LLaMA-Factory. For RL training, we customized our framework based on the VeRL training framework [^4] https://github.com/volcengine/verl.
We implement our scoring model as an API call for reward calculation during the RL training process. During our RL training, we observe that the bottleneck is the reward calculation process when using a single scoring model to evaluate rollouts from different actors. To maximize training efficiency, we deploy multiple scoring models on different GPUs in a different node (specifically, 8 scoring models in our case), and each actor calls a specific scoring model according to its local rank. This approach significantly increased our RL training efficiency.
D.2 ScoreRS Training
Table 15: Hyperparameter for training our ScoreRS